Find





lim [1/(2n+1)^2


+2/(2n+2)^2+3/(2n+3)^2+..3/25n] n→ ∞

1
Expert's answer
2022-02-27T12:10:24-0500

"\lim\limits_{n\to+\infty} [\frac{1}{(2n+1)^2}+\frac{2}{(2n+2)^2}+\frac{3}{(2n+3)^2}+..\frac{3n}{(5n)^2}]=?"

Note that

"[\frac{1}{(2n+1)^2}+\frac{2}{(2n+2)^2}+\dots+\frac{3n}{(5n)^2}]+2n[\frac{1}{(2n+1)^2}+\frac{1}{(2n+2)^2}+\dots+\frac{1}{(5n)^2}]"

"=\frac{2n+1}{(2n+1)^2}+\frac{2n+2}{(2n+2)^2}+\dots+\frac{5n}{(5n)^2}=\frac{1}{2n+1}+\frac{1}{2n+2}+\dots+\frac{1}{5n}"

Therefore,

"\frac{1}{(2n+1)^2}+\frac{2}{(2n+2)^2}+\dots+\frac{3n}{(5n)^2}=\sum\limits_{k=2n+1}^{5n}\frac{1}{k}-2n\sum\limits_{k=2n+1}^{5n}\frac{1}{k^2}" (*)

We will get an integral estimate of both terms on the right-hand side of this equality.

For all "x\in[k-1,k]" we have "\frac{1}{x+1}\leq\frac{1}{k}\leq \frac{1}{x}". Thus,

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k} \leq \sum\limits_{k=2n+1}^{5n}\int\limits_{k-1}^{k}\frac{dx}{x}=\int\limits_{2n}^{5n}\frac{dx}{x}=\log\frac{5n}{2n}=\log\frac{5}{2}" ,

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k}\geq \sum\limits_{k=2n+1}^{5n}\int\limits_{k-1}^{k}\frac{dx}{x+1}=\int\limits_{2n}^{5n}\frac{dx}{x+1}=\log\frac{5n+1}{2n+1}=\log\frac{5}{2}+O(\frac{1}{n})" ,

Therefore,

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k}=\log\frac{5}{2}+O(\frac{1}{n})"

Similarly,

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k^2} \leq \sum\limits_{k=2n+1}^{5n}\int\limits_{k-1}^{k}\frac{dx}{x^2}=\int\limits_{2n}^{5n}\frac{dx}{x^2}=\frac{1}{2n}-\frac{1}{5n}=\frac{3}{10n}"

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k^2} \geq \sum\limits_{k=2n+1}^{5n}\int\limits_{k-1}^{k}\frac{dx}{(x+1)^2}=\int\limits_{2n}^{5n}\frac{dx}{(x+1)^2}=\frac{3n}{(2n+1)(5n+1)}" "=\frac{3}{10n}+O(\frac{1}{n^2})"

Therefore,

"\sum\limits_{k=2n+1}^{5n}\frac{1}{k^2}=\frac{3}{10n}+O(\frac{1}{n^2})"

Let's apply the obtained asymptotic estimates to equality (*).

"\frac{1}{(2n+1)^2}+\frac{2}{(2n+2)^2}+\dots+\frac{3n}{(5n)^2}" "=\log\frac{5}{2}+O(\frac{1}{n})-2n(\frac{3}{10n}+O(\frac{1}{n^2}))" "=\log\frac{5}{2}-\frac{3}{5}+O(\frac{1}{n})"

"\lim\limits_{n\to+\infty} [\frac{1}{(2n+1)^2}+\frac{2}{(2n+2)^2}+\frac{3}{(2n+3)^2}+..\frac{3n}{(5n)^2}]" "=\log\frac{5}{2}-\frac{3}{5}"


Answer. "\log\frac{5}{2}-\frac{3}{5}"